# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 5.1 Single-coalition-first action models
Second, we show (3b). Let $\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, and $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{\mathrm{C}}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, where $\sigma_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma_{c_{n}}\in\mathtt{JA}_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, where $\sigma^{\prime}_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma^{\prime}_{c_{n}}\in\mathtt{JA}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $1\leq i\leq n$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{m}}(s,\sigma^{\prime}_{c_{m}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume (3). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
@@
  • Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. By (3a), $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
Second, we show (2b). We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ but $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Note $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$. Assume $\sigma_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Note $\mathrm{C}\cap\mathrm{D}=\emptyset$. Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction. Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(s)$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$ and $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}})$. Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ is a fusion of $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}$. By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction.
@@
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counter-example.
@@
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. Convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
@@ 5.2 Single-coalition-first neighborhood models
[…]
@@
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST})-\{\emptyset\})$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a general cover of $\mathtt{suc}(s)$.
@@
Definition 17 (Neighborhood functions).
@@
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathrm{C}=\emptyset\text{ and }\mathtt{suc}(s)\neq\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathrm{C}\neq\emptyset\end{cases}$
Fact 6.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
@@
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}=\emptyset$ and $\mathtt{suc}(s)\neq\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Let $t\in\mathtt{suc}(s)$. Then there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in X_{a_{1}}$, …, there is $X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in X_{a_{n}}$. Then $t\in X_{a_{1}}\cap\dots\cap X_{a_{n}}$. Note $X_{a_{1}}\cap\dots\cap X_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
@@
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if and only if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if and only if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if and only if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
As above, we let the eight strings in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
For any single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathtt{X}\in\mathrm{ES}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
@@
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Models based on so-called tight alternating transition systems are single-coalition-first neighborhood $\mathtt{SID}$-models.

### v2

@@ 4.2 Single-coalition-first action models
Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. It suffices to show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. By (2b), it suffices to show that there is $\sigma_{\overline{\mathrm{C}}}\in\mathtt{JA}_{\overline{\mathrm{C}}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Clearly, what we want holds. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there is $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})$, …, there is $\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$ such that $w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. By (2b), $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
Second, we show (3b). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$. We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$. Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}=\emptyset$. It is easy to see $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $i$ such that $1\leq i\leq n$. By (1b), $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$. By (1b), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume (3).
@@
Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which equals to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
@@
Second, we show (2b). Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Assume $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We want to get a contradiction.
We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$. Assume $\sigma_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Note $\mathrm{C}\cap\mathrm{D}=\emptyset$. Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$, which is impossible. Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$ and $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}})$. Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ is a fusion of $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}$. By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We have a contradiction.
@@
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counterexample.
@@
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. It can be verified convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
@@ 4.3 Single-coalition-first neighborhood models
Let $\mathtt{ST}$ be a nonempty set of states and $\Delta_{1},\Delta_{2}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}$. Define $\Delta_{1}\odot\Delta_{2}=\{\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\mid\mathrm{Y}_{1}\in\Delta_{1},\mathrm{Y}_{2}\in\Delta_{2},\text{and }\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset\}$. Let $I$ be a nonempty set of indices. Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected. Note $\bigodot\{\Delta\}=\Delta$.
@@
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$.
@@
Definition 17 (Neighborhood functions of single-coalition-first neighborhood models).
@@
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}=\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}\neq\emptyset\end{cases}$
Fact 4.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
@@
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Let $t\in\mathtt{suc}(s)$. We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$ for every $a\in\mathtt{AG}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$. Then $t\in\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$. Note $\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
@@
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
As above, we let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
For any $\mathtt{X}\in\mathrm{ES}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
@@
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Goranko and Jamroga [4] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.